A dental X-ray for imaging impacted wisdom teeth produces an effective dose of 10 μSv. If a dental X-ray machine emits X-rays with an energy of 6.00x10–17 J each, how many of these X-rays must be absorbed per kilogram of tissue to produce an effective dose of 10 μSv? Assume the RBE of these X-rays is 1.00.

X-rays/kg
Answer #1

Dose produced by dental X-ray = 10 μSv =10-5 Sv = 10-5 J/kg (1 Sv = 1J/kg)

Energy of X-ray = 6.00 x 10-17 J

RBE = 1.00

Number of X-rays absorbed = 10-5 J/kg /[1.00 x 6.00 x 10-17 J]

Number of X-rays absorbed = 1.667 x 1011 X-rays/kg